. NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE Profit Margin and...

.

NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE

Profit Margin and Debt Ratio

Assume you are given the following relationships for the Haslam Corporation:

Sales/total assets 1.8
Return on assets (ROA) 4%
Return on equity (ROE) 5%

Calculate Haslam's profit margin and liabilities-to-assets ratio. Do not round intermediate calculations. Round your answers to two decimal places.

a.) Profit margin:   % ?

b.) Liabilities-to-assets ratio:   % ?

c.) Suppose half of its liabilities are in the form of debt. Calculate the debt-to-assets ratio. Do not round intermediate calculations. Round your answer to two decimal places.

  %   ?
Answer #1

Answer:

a) Profit Margin=RoA*Assets/Sales=4%*1/1.8=2.22%
b) Liabilities/Assets=1-RoA/RoE=1-4%/5%=20.00%
c) Debt=0.5*Liabilities
Debt to assets ratio=0.5*Liabilities/Assets
=0.5*20%=10.00%